how long will it take a bus traveling at 60 miles per hour to overtake a car traveling at 40 mph if the car had a 1.5-hour head start?
Answer by stanbon(75887)

how long will it take a bus traveling at 60 miles per hour to overtake a car traveling at 40 mph if the car had a 1.5-hour head start?
---
slower car distance: r*t = 40x miles
---
faster car distance: r*t = 60(x-1.5) = 60x - 90 miles
----
Equation:
distance = distance
40x = 60x - 90
20x = 90
x = 9/2 = 4.5 hours
=======================
Cheers,
Stan H.